OpenStudy (jade):

Find the product of (4x − 3y)2.

OpenStudy (princeharryyy):

\[(4x-3y)^2 = (4x-3y)(4x-3y) = (4x)^2 +(3y)^2 -2\times(4x)\times(3y)\]

OpenStudy (princeharryyy):

\[=16x^2 +9y^2 -24xy\]

OpenStudy (jade):

so is this adding

OpenStudy (princeharryyy):